A tennis player receives a shot with the ball (0.0600 kg) traveling horizontally at 52.0 m/s and returns the shot with the ball traveling horizontally at 32.0 m/s in the opposite direction. (Assume the initial direction of the ball is in the -x direction.)

(a) What is the impulse delivered to the ball by the racquet?
?

(b) What work does the racquet do on the ball?

How many Joules?
Answer #1
mass of the ball m = 0.06 Kg
v1 = -52 m/s
v2 = 32 m/s
(a)
The impulse delivered to the ball = change in momentum
                                                 = m(v2 - v1)
                                                 = 0.06(32-(-52))
                                                 = 0.06(32+52)
                                                 = 0.06*84
                                                 = 5.04 Kgm/s   
(b)
work done on the ball = change in it's kinetic energy
                                 = (m/2)(v12 - v22)  
                                = 0.03*[2704 - 1024]
                                 = 50.4 J  
Answer #2
given that
      mass of the ball m = 0.0600kg
   intial velocity vi = -52 m/s
final velocity vf = 32 m/s
impulse
             FΔt = Δp
                   = m( vf - ( -vi)
                    = 5.04 N.s
(b) wor done is change in kinetic energy
               w = ΔKE
                   = ( 1/2) m ( vf2 - vi2)
                  = -50.4J
negative sign means rocket is supplying work to the ball
Answer #3

Impulse = Change in momentum.

J = Δp
J = mΔv
J = .06(-32 - 52)
J = -5.04 N*s

Work = ΔKE
= (m/2)(v12 - v22)
= 0.03*[2704 - 1024]
= 50.4 J